H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,800,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $3,010,000 in annual sales, with costs of $2,030,000. If the tax rate is 23 percent, what is the OCF for this project? (Do not round intermediate calculations and round your answer to 2 decimal places,  e.g., 32.16.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,800,000/3)=$933333.333

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(3,010,000-2,030,000)(1-0.23)+(0.23*933333.333)

=$969266.67(Approx).